Solve the following equations. 
 (3y + 4)/(2 – 6y) = -2/5 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(3y+4)/(2–6y)=-2/5
⇒3y+4=-2/5(2–6y)
⇒5(3y+4)=-2(2-6y)
⇒15y+20=-4+12y
⇒15y-12y=-4-20
⇒3y=-24
⇒y=-8

Thanks